5) What is the volume occupied by 35.2 g of methane gas (CH4) at 25°C and 1.0 atm? (3pts)

please help
Answer #1

Molar mass of CH4,

MM = 1*MM(C) + 4*MM(H)

= 1*12.01 + 4*1.008

= 16.042 g/mol

mass(CH4)= 35.2 g

use:

number of mol of CH4,

n = mass of CH4/molar mass of CH4

=(35.2 g)/(16.04 g/mol)

= 2.194 mol

Given:

P = 1.0 atm

n = 2.1942 mol

T = 25.0 oC

= (25.0+273) K

= 298 K

use:

P * V = n*R*T

1 atm * V = 2.1942 mol* 0.08206 atm.L/mol.K * 298 K

V = 53.6567 L

Answer: 53.7 L